Write the expression as the sine, cosine, or tangent of an angle.
sin 48° cos 15° - cos 48° sin 15°
And also:
Find all solutions to the equation.
cos x = sin x
And:
Verify the identity.
4 csc 2x = 2 csc2x tan x

Answer by lwsshak3(11628)

Write the expression as the sine, cosine, or tangent of an angle.
sin 48° cos 15° - cos 48° sin 15°
Identity: sin(s-t)=sin s cos t-cos s sin t
sin 48° cos 15° - cos 48° sin 15°=sin(48º-15º)=sin 33º
..
And also:
Find all solutions to the equation.
cos x = sin x
tan x=cos x/sin x=1
x=45º+360ºk, x=225º+360ºk, k=integer
..
And:
Verify the identity.
4 csc(2x) = 2 csc^2x tan x
start with right side
2csc^2x tanx
=2/sin^2x*sinx/cosx
=2/sinx cosx
multiply top and bottom by 2
=4/2sinx cosx
=4/sin(2x)
=4csc(2x)
verified: right side=left side